# The JWST Galactic Center Survey A White Paper

### v6

@@ 4.4 Astrometry
Astrometry will be done in F210M (except for highly embedded sources, which may be significantly brighter in F480M). With an angular resolution of , we can thus measure relative astrometric positions with a precision of mas for all sources with a signal-to-noise , corresponding to an observed F210M magnitude of about mag, where we expect to be still above the confusion limit. Relative precision alignment between different epochs and instruments can be obtained with several thousands of stars per image from each epoch. With relative astrometric uncertainties of mas for the stars in two given epochs, the alignment uncertainty can then be kept below 1 mas (considerations based on Schödel et al. 2009; Shahzamanian et al. 2019, 2022). With two epochs spaced five (ten) years apart, we can thus obtain proper motion measurements with an accuracy of better than (0.15) mas yr (conservative estimation based on the faintest stars) for an estimated ten million stars (extrapolated from GALACTICNUCLEUS, see section 5). This precision is well below the velocity dispersion of the NSD and inner bar and will allow us to find young star clusters and associations in the form of co-moving groups. We will also be able to measure the proper motions of deeply embedded YSOs and thus of the associated molecular clouds.
@@ 4.5 Strategy
The proposed GC survey should be carried out over various epochs to enable proper motion measurements, to decrease risk, and to facilitate the allocation of observations. Since practically all science cases require an accurate determination of interstellar extinction/reddening we consider imaging at F210M and F480M to have the highest priority and it should be carried out in the first epoch. The observations at the shortest wavelengths, where extinction is highest and source confusion therefore lowest, could potentially be carried out with WFC3/HST at F127M.
@@
A third and, possibly, fourth epoch will be observed five and ten years after the first epoch in an identical way to the first one. These epochs will serve for proper motion measurements. While epoch 2 can be used, in a limited way, for proper motions, it may be too close in time and with different filters and therefore not ideal for proper motion measurements. One of the epochs may include the F140M filter instead of F210M, to obtain the best possible wavelength leverage for CMDs.
@@
5 Synergy with existing VLT and HST surveys, proper motions
@@
The HST Paschen- survey (Wang:2010fk; Dong et al. 2011) covers an area of arcmin (35 pc) with NICMOS F187N and F190N observations. Its catalog contains point sources.
Libralato et al. (2021) present a catalog of point sources in two epochs from HST WFC3 observations, covering three disconnected fields of arcmin and one of arcmin at the GC down to magnitudes of 22 mag. They measure the proper motions of about stars with an uncertainty of mas yr.
The GALACTICNUCLEUS survey (Nogueras-Lara et al. 2019) covers an area of 0.3 deg, or about 6000 pc. It provides photometry in the J, H, and Ks bands with photometric uncertainties better than 0.1 mag (J) and mag (H and K) for over 3 million stars. At angular resolution, GALACTICNUCLEUS is the most complete survey of the GC up to date. A second epoch of H-band GALACTICNUCLEUS observations was acquired in 2021/2022 for proper motion measurements (PI Shahzamanian, ESO ID 105.20CM.001, not published yet).
The “Near Infrared Inner 600 pc Photometric Legacy Survey” (PI M. Zoccali, ESO Programme ID 0103.B-0262) used the GRAAL ground layer AO system of the wide field NIR camera HAWK-I at the VLT to image the central deg of the Milky Way at angular resolutions of in the H and Ks bands. The data have not yet been fully published.
@@
GALACTICNUCLEUS covers about 70% of the field proposed for the JWST NIRCam survey. There are two significant synergies between GALACTICNUCLEUS and the proposed NIRCam observations. (1) GALACTICNUCLEUS provides photometry at wavelengths shorter than m. In combination with the de-reddening of individual stars provided by the proposed F210M and F480M photometry, this will mean improved stellar classification (intrinsic stellar color differences for different types of stars reach up to 0.7 mag in , see Fig. 4 ). (2) There exist already two epochs of GALACTICNUCLEUS (2015, 2021). Combined with a single epoch of NIRCam imaging, we can obtain precision proper motions with an uncertainty mas yr for about stars, as described below.
Proper motions. The velocity dispersion of Bulge stars perpendicular to the Galactic Plane (and thus to the NSD) is mas/yr. The velocity dispersion of the NSD is significantly smaller mas/yr (Shahzamanian et al. 2022). The internal velocity dispersion of clusters is mas yr (Rui et al. 2019; Hosek et al. 2019). In order to be able to separate stellar populations and identify co-moving groups, we therefore aim at a proper motion uncertainty of better than mas yr (see also Matsunaga:2018dk, , who argue for a similar precision).
We discuss proper motion measurements using the example of GALACTICNUCLEUS here. GALACTICNUCLEUS is ideal for proper motion measurements, because of its large area and high angular resolution, and because the data reduction pipeline minimises residual differential atmospheric tip-tilt effects. Relative astrometric uncertainties of mas are obtained in GALACTICNUCLEUS for stars of mag ( is the preferred band for astrometry in GALACTICNUCLEUS, see Shahzamanian et al. 2022; Martinez-Arranz:2022wj). Assuming a mean extinction of mag, mag corresponds to [F210M]16.5 (17.5) mag, fainter than the Red Clump at the distance and extinction of the GC. Shahzamanian et al. (2022) presented a pilot proper motion study for the area common to GALACTICNUCLEUS and the HST Paschen- survey. Major limitations were the relatively low SNR and small field-of-view of the NICMOS data. In particular the latter increased the uncertainty of epoch alignment. When combining NIRCam observations with GALACTICNUCLEUS, these problems will not be important. Due to the sensitivity and large field-of-view of NIRCam, we will be able to use on the order one thousand star for precise alignment between the data sets, thus limiting alignment uncertainties to mas.
@@
With estimates based on Sgr C observations, the survey area, and the SFH, we arrive at a number of 10 million stars for which proper motions could be measured with a second epoch of NIRCam imaging. The uncertainties of stars above the confusion limit () would be mas yr, assuming 1 mas relative astrometry, 1 mas alignment uncertainty etween epochs and a five year time baseline. Such revolutionary data would enable exploring the GC kinematically down to almost solar mass main sequence stars.
@@ Appendix A Appendix: Detection & Classification of YSOs
To improve the realism of the YSO modeling, we apply a set of selections to the Robitaille (2017) model grid that excludes unrealistic sources and classifies YSO models by their theoretical stage, allowing some distinction between their ages. Figure 8 shows a simulated population of YSOs in the selected F212N and F480M filters. The model population is generated assuming a constant star formation rate of 0.1 yr, timescales of the Stage 0, I, and II evolutionary stages of 0.1, 0.5, and 5.0 Myr, respectively, and an assumed luminosity scaling. They are sampled from a Salpeter IMF. Only stars with M () are included, as these are roughly the YSOs expected to be dominated by stellar luminosity rather than accretion luminosity. For lower-mass sources, the luminosity scatters between extremely faint () and quite bright () depending on their accretion rate (Dunham et al. 2014). This is a very coarse population synthesis, but it suggests that there are detectable M Stage 0/I YSOs across the CMZ, most of which are Stage I (i.e., they have envelopes). There should be Stage 0 and Stage I sources from 0.5-4 M, this time adopting a Kroupa mass function. An unknown fraction of these sources may be detectable along with the above M YSOs if they are accreting at M yr. There may be about as many Stage II (disk, but no envelope) objects that may have substantial infrared excess.
@@ B.1 Shorter wavelengths
Shorter wavelengths, such as the near-infrared J and H bands, are highly desirable for stellar classification. They are more useful than the selected K-band (F210M) for determining stellar properties. Additionally, the higher angular resolution results in higher accuracy astrometry.

### v7

@@ 4.4 Astrometry
Astrometry will be done in F212N (except for highly embedded sources, which may be significantly brighter in F480M). With an angular resolution of , we can thus measure relative astrometric positions with a precision of mas for all sources with a signal-to-noise , corresponding to an observed F212N magnitude of about mag, where we expect to be well above the confusion limit. Relative precision alignment between different epochs and instruments can be obtained with several thousands of stars per image from each epoch. With relative astrometric uncertainties of mas for the stars in two given epochs, the alignment uncertainty can then be kept below 1 mas (considerations based on Schödel et al. 2009; Shahzamanian et al. 2019, 2022). With two epochs spaced five (ten) years apart, we can thus obtain proper motion measurements with an accuracy of better than (0.15) mas yr-1 (conservative estimation based on the faintest stars) for an estimated ten million stars (extrapolated from GALACTICNUCLEUS, see section 5). This precision is well below the velocity dispersion of the NSD and inner bar and will allow us to find young star clusters and associations in the form of co-moving groups. We will also be able to measure the proper motions of deeply embedded YSOs and thus of the associated molecular clouds.
@@ 4.5 Strategy
The proposed GC survey should be carried out over various epochs to enable proper motion measurements, to decrease risk, and to facilitate the allocation of observations. Since practically all science cases require an accurate determination of interstellar extinction/reddening we consider imaging at F212N and F480M to have the highest priority and it should be carried out in the first epoch. The observations at the shortest wavelengths, where extinction is highest and source confusion therefore lowest, could potentially be carried out with WFC3/HST at F127M.
@@
A third and, possibly, fourth epoch will be observed five and ten years after the first epoch in an identical way to the first one. These epochs will serve for proper motion measurements. While epoch 2 can be used, in a limited way, for proper motions, it may be too close in time and with different filters and therefore not ideal for proper motion measurements. One of the epochs may include the F140M filter instead of F212N, to obtain the best possible wavelength leverage for CMDs.
@@
5 Synergy with existing VLT, HST, and Roman surveys, proper motions
@@
The HST Paschen- survey (Wang:2010fk; Dong et al. 2011) covers an area of arcmin2 (35 pc2) with NICMOS F187N and F190N observations. Its catalog contains point sources.
Libralato et al. (2021) present a catalog of point sources in two epochs from HST WFC3 observations, covering three disconnected fields of arcmin2 and one of arcmin2 at the GC down to magnitudes of 22 mag. They measure the proper motions of about stars with an uncertainty of mas yr-1.
The GALACTICNUCLEUS survey (Nogueras-Lara et al. 2019) covers an area of 0.3 deg2, or about 6000 pc2. It provides photometry in the J, H, and Ks bands with photometric uncertainties better than 0.1 mag (J) and mag (H and Ks) for over 3 million stars. At angular resolution, GALACTICNUCLEUS is the most complete survey of the GC up to date. A second epoch of H-band GALACTICNUCLEUS observations was acquired in 2021/2022 for proper motion measurements (PI Shahzamanian, ESO ID 105.20CM.001, not published yet).
The “Near Infrared Inner 600 pc Photometric Legacy Survey” (PI M. Zoccali, ESO Programme ID 0103.B-0262) used the GRAAL ground layer AO system of the wide field NIR camera HAWK-I at the VLT to image the central deg2 of the Milky Way at angular resolutions of in the H and Ks bands. The data have not yet been fully published.
@@
GALACTICNUCLEUS covers about 70% of the field proposed for the JWST NIRCam survey. There are two significant synergies between GALACTICNUCLEUS and the proposed NIRCam observations. (1) GALACTICNUCLEUS provides photometry at wavelengths shorter than m. In combination with the de-reddening of individual stars provided by the proposed F212N and F480M photometry, this will mean improved stellar classification (intrinsic stellar color differences for different types of stars reach up to 0.7 mag in , see Fig. 4 ). (2) There exist already two epochs of GALACTICNUCLEUS (2015, 2021). Combined with a single epoch of NIRCam imaging, we can obtain precision proper motions with an uncertainty mas yr-1 for about stars, as described below.
The Roman Space Telescope Galactic Plane Survey will image the GC with the F129, F158, F184, and F213 filters (R. Benjamin, priv.comm.). The GC NIRCam observations proposed in this White Paper will provide the community with essential complementary data to break degeneracies between stellar colours and extinction in these bands (as illustrated in Fig. 4; changing NIRCam F212N/F140M to WFI F213/129 will result in no change in the CMDs) and to identify closely blended stars in the lower resolution Roman images.
Proper motions. The velocity dispersion of Bulge stars perpendicular to the Galactic Plane (and thus to the NSD) is mas/yr. The velocity dispersion of the NSD is significantly smaller mas/yr (Shahzamanian et al. 2022). The internal velocity dispersion of clusters is mas yr (Rui et al. 2019; Hosek et al. 2019). In order to be able to separate stellar populations and identify co-moving groups, we therefore aim at a proper motion uncertainty of better than mas yr-1 (see also Matsunaga:2018dk, , who argue for a similar precision).
We discuss proper motion measurements using the example of GALACTICNUCLEUS here. GALACTICNUCLEUS is ideal for proper motion measurements, because of its large area and high angular resolution, and because the data reduction pipeline minimises residual differential atmospheric tip-tilt effects. Relative astrometric uncertainties of mas are obtained in GALACTICNUCLEUS for stars of mag ( is the preferred band for astrometry in GALACTICNUCLEUS, see Shahzamanian et al. 2022; Martinez-Arranz:2022wj). Assuming a mean extinction of mag, mag corresponds to [F212N]16.5 (17.5) mag, fainter than the Red Clump at the distance and extinction of the GC. Shahzamanian et al. (2022) presented a pilot proper motion study for the area common to GALACTICNUCLEUS and the HST Paschen- survey. Major limitations were the relatively low SNR and small field-of-view of the NICMOS data. In particular the latter increased the uncertainty of epoch alignment. When combining NIRCam observations with GALACTICNUCLEUS, these problems will not be important. Due to the sensitivity and large field-of-view of NIRCam, we will be able to use on the order one thousand star for precise alignment between the data sets, thus limiting alignment uncertainties to mas.
@@
With estimates based on Sgr C observations, the survey area, and the SFH, we arrive at a number of 10 million stars for which proper motions could be measured with a second epoch of NIRCam imaging. The uncertainties of stars above the confusion limit () would be mas yr-1, assuming 1 mas relative astrometry, 1 mas alignment uncertainty etween epochs and a five year time baseline. Such revolutionary data would enable exploring the GC kinematically down to almost solar mass main sequence stars.
@@ Appendix A Appendix: Detection & Classification of YSOs
To improve the realism of the YSO modeling, we apply a set of selections to the Robitaille (2017) model grid that excludes unrealistic sources and classifies YSO models by their theoretical stage, allowing some distinction between their ages. Figure 8 shows a simulated population of YSOs in the selected F212N and F480M filters. The model population is generated assuming a constant star formation rate of 0.1 yr-1, timescales of the Stage 0, I, and II evolutionary stages of 0.1, 0.5, and 5.0 Myr, respectively, and an assumed luminosity scaling. They are sampled from a Salpeter IMF. Only stars with M⊙ () are included, as these are roughly the YSOs expected to be dominated by stellar luminosity rather than accretion luminosity. For lower-mass sources, the luminosity scatters between extremely faint () and quite bright () depending on their accretion rate (Dunham et al. 2014). This is a very coarse population synthesis, but it suggests that there are detectable M⊙ Stage 0/I YSOs across the CMZ, most of which are Stage I (i.e., they have envelopes). There should be Stage 0 and Stage I sources from 0.5-4 M⊙, this time adopting a Kroupa mass function. An unknown fraction of these sources may be detectable along with the above M⊙ YSOs if they are accreting at M⊙ yr-1. There may be about as many Stage II (disk, but no envelope) objects that may have substantial infrared excess.
@@ B.1 Shorter wavelengths
Shorter wavelengths, such as the near-infrared J and H bands, are highly desirable for stellar classification. They are more useful than the selected K-band (F212N) for determining stellar properties. Additionally, the higher angular resolution results in higher accuracy astrometry.